Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

UCD Air Rifle Open

Distribution

Dr. Trevor West DUCAC Chairman
Aubrey Storey DUCAC Secretary
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record
Susan Thompson Trinity News

Summary

Venue UCD Rifle Club, Belfield
Date 28-29 October 2001
Discipline 10m Air Rifle

DURC Attendees

Steven Watterson  Captain
Mark Dennehy  Secretary
Ben Jones  Treasurer
Peter Byrne  Adrian Costigan
Nithin Santhosh  Dave Barron
Colman Marcus-Quinn  Ljiljana Skrba

Daniel O’Neill

Scores

Steven Watterson 571 (669.8)
Mark Dennehy 550
Ben Jones 491
Peter Byrne 475
Adrian Costigan 474
Nithin Santosh 430
Dave Banon 377
Colman Marcus-Quinn 359
Ljiljana Skrba 321
Daniel O’Neill 281

Prizes Won

Steven Watterson Third Place Overall

Third Place, Class A

Mark Dennehy Third Place, Class B
Peter Byrne Second Place, Class D
Adrian Costigan Third Place, Class D

Report

DURC sent a moderate contingent to the UCD Air Rifle Open this weekend. This is the first Air Rifle competition of the season that we can get large numbers of people to, owing to the nearby location of UCD. We had ten people attend in total, breaking down to five experienced shooters and five novices, a reasonable demographic.

Steven, Mark, Peter and Adrian all shot personal bests, and secured prizes. Mark, at 550, is now set to enter class A after the next competition. Nithin, Dave, Colman and Ljiljana all shot very well for their first competition, especially considering that they had only had two or three half-hour details of shooting prior to the competition.

The next competition will be the Co. Wicklow Open in Rathdrum on November 4, and DURC will be sending another squad to attend.

Peter shooting opposite Collette Kane of Wilkinstown Target Shooting Club; and Ben shooting opposite Robert Clarke of East Antrim Rifle Club

Misé le meas,

Mark Dennehy, DURC Secretary